# Unfitted finite element interpolated neural networks

### v1

@@ 4.1.1. The effect of Nitsche coefficient for Poisson equation
In FEM, the choice of Nitsche coefficient $\gamma$ significantly affects solution accuracy. Thus, we first investigate how different $\gamma$ values impact the training of FEINNs for solving the Poisson equation. The physical domain is a disk centred at $(0.5,0.5)$ with a radius of $R=0.4$, and the background mesh is a $50\times 50$ uniform grid over $[0,1]^{2}$. The trial FE space has an order of $k_{U}=3$. We use different $\gamma$ values in the loss function and train NNs starting from the same initial parameters for each $\gamma$. We adopt the following smooth solution from [Badia2024]:
@@
Fig. 3 shows the $L^{2}$ and $H^{1}$ error histories for the interpolated NNs and the NNs themselves during training with different values of $\gamma$ in the loss function. An immediate observation is that a large $\gamma$ value (e.g., $\gamma=100$) leads to a much slower convergence. However, convergence accelerates when $\gamma$ is small, and the corresponding FEINNs eventually converge to the $\left\lVert u-\pi_{h}(u)\right\rVert$ lines. Another important finding is that, consistent with observations in [Badia2024, Badia2024adaptive], the NNs themselves can outperform their FE interpolations in terms of $L^{2}$ and $H^{1}$ errors. As shown in the second column of Fig. 3, the FEINN curves converge to the $\left\lVert u-\pi_{h}(u)\right\rVert$ lines, while the NN curves eventually fall below these lines when $\gamma$ is suitable.
@@
Among the tested values, $\gamma=0.01$ achieves the fastest convergence, so for convenience, we simply select $\gamma=0.01$ for the remaining experiments in this work. However, we emphasise that unfitted FEINNs are robust to variations in $\gamma$ across a wide range. As shown in Fig. 3, the differences among convergence curves for $\gamma\leq 10$ values are not significant. For instance, $\gamma=0.0001$ and $\gamma=10$ yield similar error histories. Therefore, as long as $\gamma$ is “small enough” we expect unfitted FEINNs to perform well.
@@ 4.1.2. Moving domain test for the Poisson equation
We now examine the robustness of unfitted FEINNs to changes in the relative position between the physical domain and the background mesh. We adapt the moving disk experiment from [Badia2018, Sect. 6.2], where a disk with radius $R=0.19$ moves along the anti-diagonal of the background box $[0,1]^{2}$ by shifting its centre $(x,x)$. As $x$ varies, the relative position of the disk with respect to the background mesh changes, leading to a comprehensive exploration of different cut cell configurations. The goal of the experiment is to study how the accuracy of the trained NNs might be affected by small cut cells. We use $k_{U}=3$ and a uniform background mesh with $50\times 50$ quadrilaterals. The value of $x$ ranges from 0.2 to 0.8, with a small step size of 0.005 to increase the probability of encountering small cut cells. The true solution is still given by (11).
Fig. 4 displays the $L^{2}$ and $H^{1}$ errors of the identified solution versus the centre position of the moving disk. The NNs start with the same initial parameters for each $x$ value. For the interpolated NNs, both $L^{2}$ and $H^{1}$ errors closely follow the flat $\left\lVert u-\pi_{h}(u)\right\rVert$ lines. While there are fluctuations in the NN errors, they consistently remain below the $\left\lVert u-\pi_{h}(u)\right\rVert$ lines. Notably, the $H^{1}$ errors of the NNs are steadily more than an order of magnitude lower than the FEINN errors. These observations highlight that the unfitted FEINN method is highly robust to small cut cells without using any kind of FE stabilisation such as AgFEM or CutFEM to compute the residual. In the context of FEINNs, such stabilisations are required only when using the preconditioned loss function (9), where the discrete dual norms of the residual must be computed.
@@ 4.1.3. Convergence tests for the Poisson equation
In previous experiments, we have established that a Nitsche coefficient of $\gamma=0.01$ yields accurate solutions for $k_{U}\in\{2,3\}$ and confirmed the robustness of unfitted FEINNs against small cut cells. We now confidently proceed to investigate the convergence properties of FEINNs with respect to the trial space background mesh size ($h$) and order ($k_{U}$). We use the disk geometry centred at $(0.5,0.5)$ with radius $R=0.4$ and a uniform Cartesian background mesh on $[0,1]^{2}$. The true solution is again (11).
We first fix $k_{U}$ and perform an $h$-refinement experiment for unfitted FEINNs. The convergence results are shown in Fig. 5. Different colours distinguish between $k_{U}=2$ and $k_{U}=3$, while markers are used to differentiate between NNs and their interpolations. The dashed lines labelled as “FE interp” represent errors of $\pi_{h}(u)$ versus mesh size.
@@
All the $\left\lVert u-\pi_{h}(u)\right\rVert$ lines in Fig. 5 exhibit the expected slopes, which validates the accuracy of the analytic solution FE interpolation and confirms that the error computations use sufficient Gaussian quadrature degree. The FEINN errors always align with the $\left\lVert u-\pi_{h}(u)\right\rVert$ lines for both $k_{U}=2$ and $k_{U}=3$, indicating that FEINNs converge at the optimal rate. Both $L^{2}$ and $H^{1}$ errors for NNs are below the corresponding $\left\lVert u-\pi_{h}(u)\right\rVert$ lines, manifesting NNs’ superior accuracy compared to their interpolation counterparts and $\pi_{h}(u)$. Remarkably, the finest mesh for $k_{U}=3$ yields $H^{1}$ errors that are about two orders of magnitude lower than the FEINN errors. Additionally, the impact of NN initialisation is minimal, as the NN errors from different initialisations are closely clustered.
Given the smoothness of the analytic solution (11), we also investigate how $k_{U}$ affects the convergence rate of unfitted FEINNs through $p$-refinement tests. We fix the background mesh size and increase $k_{U}$ from 1 to 5. The experiment results are displayed in Fig. 6. We consider two mesh sizes, $h=1/14$ and $h=1/21$, differentiated by colours in the figure. Similar to the $h$-refinement test, the FEINN errors agree with the true solution interpolation errors, indicating that FEINNs accomplish the expected convergence rate as $k_{U}$ increases for both mesh sizes. The NN errors are positioned below the $\left\lVert u-\pi_{h}(u)\right\rVert$ lines, demonstrating the improved NN accuracy without interpolations. Additionally, Fig. 6 validates the choice of $\gamma=0.01$ for $k_{U}$ values ranging from 1 to 5, building on the previous subsections where $k_{U}\in\{2,3\}$.
@@
4.1.4. The effect of preconditioners on the Poisson equation
In this experiment, we explore the impact of discrete dual norms (i.e., preconditioned losses) in the training of unfitted FEINNs for the Poisson equation. We use the same disk geometry as in Sect. 4.1.3 and set $k_{U}=3$. The background box $[0,1]^{2}$ is partitioned into $50\times 50$ quadrilaterals. To increase the difficulty of the problem, we use the following analytic solution, which features a mild sharp gradient around the centre of the disk:
@@ 4.1.4. The effect of preconditioners on the Poisson equation
As described in Sect. 2.4, we consider two linearised test FE spaces: the standard active space $V_{h}^{\rm{std}}$ consisting of all internal and cut elements, and the aggregated active space $V_{h}^{\rm{ag}}$ For simplicity, we just explore as a preconditioner the action of the inverse of the discretisation matrix built from the bilinear form using the linearised test space as both trial and test spaces. However, other types of preconditioners can also be considered. For example, in [Badia2024, Sect. 6.1.3], we demonstrated that a single cycle of GMG preconditioning was effective for the Poisson equation. Since small cut cells in $V_{h}^{\rm{std}}$ may cause its discretisation matrix to become singular in the presence of rounding errors, additional ghost penalty terms (5) are required to ensure the preconditioner is well-posed. Note that in the context of unfitted FEINNs, even though the trial space may be of higher order, $V_{h}^{\rm{std}}$ is always linear. Therefore, a first-order penalty term is sufficient to stabilise the preconditioner, i.e., $k_{U}=1$ in (5). The implementation of unfitted FEINNs does not actually need to deal with high order derivatives in the case of CutFEM stabilisation, nor specialised discrete extensions required for high-order aggregated FE spaces [Badia2022c]. Following the recommendation in [Burman2015], we set $\gamma_{g}=0.1$ in this experiment. On the other hand, the $V_{h}^{\rm{ag}}$ space is designed to overcome stability issues by construction, thus no additional penalty terms are needed in the weak formulation.
In terms of computational cost, we note that the preconditioning matrix is symmetric positive definite and is built and factorised using sparse Cholesky factorisation only once before training. Since the problem is linear, this factorisation can be stored and reused throughout training. At each iteration, the inverse of the preconditioner is applied to the residual vector by forward and backward substitutions using the sparse Cholesky factor, which makes the training computationally efficient.
@@
As shown earlier in Fig. 3, large Nitsche coefficients can slow down the training convergence, so we include both the suitable $\gamma=0.01$ and a large $\gamma=100$ in the experiment to test how effectively preconditioning handles unsuitable Nitsche coefficients. For clarity in the plots, in this experiment, we just consider the $H^{1}$ primal norm in the preconditioned loss (9). However, other primal norms can also be considered. For example, in [Badia2024adaptive, Sect. 5.1.2], we showed that four different primal norms were effective for the Poisson equation.
@@
We plot the error histories for FEINNs and NNs preconditioned by different combinations of test spaces and norms in Fig. 7. For reference, we also include the unpreconditioned error histories and label them with the postfix “$V_{h}^{\rm{std}}$, NA”. We observe that without preconditioning, a large Nitsche coefficient also impedes the convergence of both FEINNs and NNs for this problem (with a solution featuring a sharp gradient around the disk centre). However, the use of dual discrete norms overcomes this issue and accelerates the convergence significantly. For instance, with $\gamma=100$, the preconditioned FEINN $L^{2}$ curve reaches $10^{-5}$ at around 800 iterations, while the unpreconditioned curve requires more than 4,000 iterations to achieve the same error level. Besides, preconditioners are also effective for the suitable Nitsche coefficient ($\gamma=0.01$), as they not only expedite the convergence of FEINNs and NNs but also enhance the NN solution accuracy. As shown in Fig. 7, the preconditioned error curves decrease more rapidly than the unpreconditioned ones. In addition, the preconditioned NN error curves in the second column of Fig. 7 drop below the unpreconditioned curves shortly after training begins.
@@
It is clear from Fig. 7 that all the preconditioned error curves for different $\gamma$ and test space combinations are very close. This implies that preconditoners are robust to variations in $\gamma$ values and that the choice of test space has only a minor impact on their effectiveness. Given its discretisation matrix requiring no additional stabilisation nor parameter tuning, we recommend using $V_{h}^{\rm{ag}}$ as the test space for preconditioners in practice. However, we reiterate that the unfitted FEINN framework only requires a first-order ghost penalty term to stabilise $V_{h}^{\rm{std}}$, which makes $V_{h}^{\rm{std}}$ also a practical choice for preconditioning.

### v2

@@ 4.1.1. The effect of Nitsche coefficient for Poisson equation
In FEM, the choice of Nitsche coefficient $\gamma$ significantly affects solution accuracy. Thus, we first investigate how different $\gamma$ values impact the training of $\ell^{2}$-FEINNs for solving the Poisson equation. The physical domain is a disk centred at $(0.5,0.5)$ with a radius of $R=0.4$, and the background mesh is a $50\times 50$ uniform grid over $[0,1]^{2}$. The trial FE space has an order of $k_{U}=3$. We use different $\gamma$ values in the loss function and train NNs starting from the same initial parameters for each $\gamma$. We adopt the following smooth solution from [Badia2024]:
@@
Fig. 3 shows the $L^{2}$ and $H^{1}$ error histories for the interpolated NNs and the NNs themselves during training with different values of $\gamma$ in the loss function. An immediate observation is that a large $\gamma$ value (e.g., $\gamma=10^{2}$) leads to a much slower convergence. However, convergence accelerates when $\gamma$ is small, and the corresponding $\pi_{h}(u_{\mathcal{N}})$ solutions eventually converge to the $\left\lVert u-\pi_{h}(u)\right\rVert$ lines. Another important finding is that, consistent with observations in [Badia2024, Badia2025adaptive], the NNs themselves can outperform their FE interpolations in terms of $L^{2}$ and $H^{1}$ errors. As shown in the second column of Fig. 3, the curves for $\pi_{h}(u_{\mathcal{N}})$ converge to the $\left\lVert u-\pi_{h}(u)\right\rVert$ lines, while the $u_{\mathcal{N}}$ curves eventually fall below these lines when $\gamma$ is suitable.
@@
Among the tested values, $\gamma=10^{-2}$ achieves the fastest convergence, so for convenience, we simply select $\gamma=10^{-2}$ for the remaining experiments in this work. However, we emphasise that $\ell^{2}$-FEINNs are robust to variations in $\gamma$ across a wide range. As shown in Fig. 3, the differences among convergence curves for $\gamma\leq 10$ values are not significant. For instance, $\gamma=10^{-4}$ and $\gamma=10$ yield similar error histories. Therefore, as long as $\gamma$ is “small enough” we expect $\ell^{2}$-FEINNs to perform well.
4.1.2. Comparison of PINNs and unfitted FEINNs
This experiment compares the computational efficiency of unfitted FEINNs with PINNs. The loss function for PINNs [Raissi2019] applied to the Poisson equation (1) reads:
$\mathscr{L}(u_{\mathcal{N}})=\frac{1}{N_{\rm int}}\sum_{i=1}^{N_{\rm int}}\left[\boldsymbol{\Delta}u_{\mathcal{N}}(x_{i},y_{i})+f(x_{i},y_{i})\right]^{2}+\tau_{b}\frac{1}{N_{\rm b}}\sum_{i=1}^{N_{\rm b}}\left[g(x_{i},y_{i})-u_{\mathcal{N}}(x_{i},y_{i})\right]^{2},$
where $N_{\rm int}$ and $N_{\rm b}$ are the number of interior and boundary collocation points, respectively, and $\tau_{b}$ is a penalty coefficient. The first term in the loss imposes the PDE, while the second term weakly enforces the Dirichlet boundary condition through penalisation.
We adopt the same disk geometry and smooth analytic solution as in Sect. 4.1.1. For the $\ell^{2}$-FEINN experiments, the background mesh consists of $25\times 25$ quadrilaterals, with $k_{U}=4$ and $\gamma=10^{-2}$. To ensure a fair comparison, we adopt the strategy in [Berrone2022], where PINNs use the same number of collocation points as the number of interpolation nodes employed by FEINNs. In particular, we extract the FEINN interpolation nodes and use those inside $\Omega$ as interior collocation points for PINNs. Using the level-set function, we then uniformly generate boundary collocation points such that their total number equals to the number the interpolation nodes located outside $\Omega$ in FEINNs. In this setup, $N_{\rm int}=5013$ and $N_{\rm b}=868$. This approach ensures both methods are trained on the same interior points and the same amount of total points. However, the computational cost is more favourable to PINNs, because as opposed to Nitsche’s method in FEINNs, they do not require computing spatial derivatives on the boundary (approximately 15% of the total points).
As suggested in [Berrone2022], convergence in PINN training can be accelerated by initially optimising with Adam and then switching to BFGS. However, convergence curves in [Berrone2022, Fig. 5] show that BFGS yields a steeper descent than Adam. Based on this observation and our preliminary experimental results, we select only BFGS for PINNs training. For the penalty coefficient $\tau_{b}$, we tested several values ranging from $10^{-4}$ to $10^{4}$ and found that $\tau_{b}=10^{3}$ yielded the best results for this problem. We implement PINNs in the TensorFlow framework [tensorflow2015]. For a fair comparison, $\ell^{2}$-FEINNs are also implemented and trained in TensorFlow.111 We are aware that PINNs can achieve higher accuracy with alternative optimisers, such as self-scaled BFGS and Broyden methods proposed in [Urban2025]. We also observe that different implementations of the same optimiser can yield different convergence behaviours. For instance, while SciPy’s BFGS can yield higher accuracy for PINNs with smaller NN architectures, it becomes too inefficient to use for larger NNs as the ones used in this work. A detailed exploration of these optimisers and the subtle differences in their implementations is beyond the scope of this work. All experiments are conducted on a single NVIDIA V100 GPU with 32 GB of memory. The $H^{1}$ error convergence history of NNs trained with PINNs and $\ell^{2}$-FEINNs is shown in Fig. 4.
Figure 4. Convergence history of the NN $H^{1}$ error over training iterations and wall-clock time in seconds for the 2D Poisson problem using PINNs and $\ell^{2}$-FEINNs. Both methods are implemented in TensorFlow, employ identical NN architectures with the same initialisation, and are trained on GPU using TensorFlow’s BFGS with an equal number of training points.
Let us first focus on the left subfigure of Fig. 4, which shows the convergence history of the NN $H^{1}$ error over BFGS iterations. When the penalty coefficient $\tau_{b}$ is appropriately chosen (e.g., $\tau_{b}=10^{3}$), PINNs achieve accuracy comparable to $\ell^{2}$-FEINNs in terms of $H^{1}$ error. However, the choice of $\tau_{b}$ significantly affects PINNs’ accuracy: the curve for $\tau_{b}=10^{-1}$ is placed well above that for $\tau_{b}=10^{3}$, indicating that an unsuitable $\tau_{b}$ can severely hinder PINN convergence. In contrast, as shown in Fig. 3, NNs trained by $\ell^{2}$-FEINNs converge below the $\left\lVert u-\pi_{h}(u)\right\rVert$ reference lines even with suboptimal $\gamma$ values, demonstrating unfitted FEINNs’ greater robustness to hyperparameter variation.
The right subfigure of Fig. 4 compares the computational time for training PINNs and $\ell^{2}$-FEINNs. Notably, the $\ell^{2}$-FEINN method completes 5,000 BFGS iterations in around 750 seconds, while PINNs requires around 1,400 seconds. This substantial efficiency gain of $\ell^{2}$-FEINNs stems not only from avoiding computation of spatial derivatives of NNs to accelerate loss and gradient evaluations, but also from a better-posed optimisation problem that allows BFGS to converge with fewer gradient calculations. Specifically, $\ell^{2}$-FEINNs require 11,308 gradient computations, compared to 13,031 for PINNs with $\tau_{b}=10^{3}$. Moreover, each gradient computation takes 0.0645 seconds for $\ell^{2}$-FEINNs, versus 0.1064 seconds for PINNs. These results align with the findings in [Berrone2022], where IVPINNs were also shown to be faster than PINNs and VPINNs in training time.
In summary, for smooth solutions, PINNs have the potential to achieve similar accuracy as unfitted $\ell^{2}$-FEINNs; however, their performance is sensitive to the choice of the hyperparameter $\tau_{b}$. Besides, by avoiding the computation of spatial and nested derivatives of NNs, $\ell^{2}$-FEINNs are faster. It is important to mention that, in this experiment, the level-set function enables easy generation of collocation points for PINNs. However, for more complex geometries, such as those defined by STL meshes, as discussed in Sect. 4.2.2, generating collocation points becomes challenging. In contrast, $\ell^{2}$-FEINNs can handle such complex geometries without additional effort, as demonstrated in the 3D experiments.
@@ 4.1.3. Moving domain test for the Poisson equation
We now examine the robustness of unfitted $\ell^{2}$-FEINNs to changes in the relative position between the physical domain and the background mesh. We adapt the moving disk experiment from [Badia2018, Sect. 6.2], where a disk with radius $R=0.19$ moves along the anti-diagonal of the background box $[0,1]^{2}$ by shifting its centre $(x,x)$. As $x$ varies, the relative position of the disk with respect to the background mesh changes, leading to a comprehensive exploration of different cut cell configurations. The goal of the experiment is to study how the accuracy of the trained NNs might be affected by small cut cells. We use $k_{U}=3$ and a uniform background mesh with $50\times 50$ quadrilaterals. The value of $x$ ranges from 0.2 to 0.8, with a small step size of 0.005 to increase the probability of encountering small cut cells. The true solution is still given by (11).
Fig. 5 displays the $L^{2}$ and $H^{1}$ errors of the identified solution versus the centre position of the moving disk. The NNs start with the same initial parameters for each $x$ value. For the interpolated NNs, both $L^{2}$ and $H^{1}$ errors closely follow the flat $\left\lVert u-\pi_{h}(u)\right\rVert$ lines. While there are fluctuations in the NN errors, they consistently remain below the $\left\lVert u-\pi_{h}(u)\right\rVert$ lines. Notably, the $H^{1}$ errors of the NNs are steadily more than an order of magnitude lower than the $\pi_{h}(u_{\mathcal{N}})$ errors. These observations highlight that the unfitted $\ell^{2}$-FEINN method is highly robust to small cut cells without using any kind of FE stabilisation such as AgFEM or CutFEM to compute the residual. In the context of unfitted FEINNs, such stabilisations are required only when using the discrete dual norm in the loss function (9), where the discrete dual norms of the residual must be computed.
@@ 4.1.4. Convergence tests for the Poisson equation
In previous experiments, we have established that a Nitsche coefficient of $\gamma=10^{-2}$ yields accurate solutions for $k_{U}\in\{2,3\}$ and confirmed the robustness of unfitted $\ell^{2}$-FEINNs against small cut cells. We now confidently proceed to investigate the convergence properties of $\ell^{2}$-FEINNs with respect to the trial space background mesh size ($h$) and order ($k_{U}$). We use the disk geometry centred at $(0.5,0.5)$ with radius $R=0.4$ and a uniform Cartesian background mesh on $[0,1]^{2}$. The true solution is again (11).
We first fix $k_{U}$ and perform an $h$-refinement experiment for unfitted $\ell^{2}$-FEINNs. The convergence results are shown in Fig. 6. Different colours distinguish between $k_{U}=2$ and $k_{U}=3$, while markers are used to differentiate between NNs and their interpolations. The dashed lines represent errors of $\pi_{h}(u)$ versus mesh size.
@@
All the $\left\lVert u-\pi_{h}(u)\right\rVert$ lines in Fig. 6 exhibit the expected slopes, which validates the accuracy of the analytic solution FE interpolation and confirms that the error computations use sufficient Gaussian quadrature degree. The $\pi_{h}(u_{\mathcal{N}})$ errors always align with the $\left\lVert u-\pi_{h}(u)\right\rVert$ lines for both $k_{U}=2$ and $k_{U}=3$, indicating that $\ell^{2}$-FEINNs converge at the optimal rate. Both $L^{2}$ and $H^{1}$ errors for NNs are below the corresponding $\left\lVert u-\pi_{h}(u)\right\rVert$ lines, manifesting NNs’ superior accuracy compared to their interpolation counterparts and $\pi_{h}(u)$. Remarkably, the finest mesh for $k_{U}=3$ yields $H^{1}$ errors that are about two orders of magnitude lower than the $\pi_{h}(u_{\mathcal{N}})$ errors. Additionally, the impact of NN initialisation is minimal, as the NN errors from different initialisations are closely clustered.
Given the smoothness of the analytic solution (11), we also investigate how $k_{U}$ affects the convergence rate of unfitted $\ell^{2}$-FEINNs through $p$-refinement tests. We fix the background mesh size and increase $k_{U}$ from 1 to 5. The experiment results are displayed in Fig. 7. We consider two mesh sizes, $h=1/14$ and $h=1/21$, differentiated by colours in the figure. Similar to the $h$-refinement test, the $\pi_{h}(u_{\mathcal{N}})$ errors agree with the true solution interpolation errors, indicating that the method accomplishes the expected convergence rate as $k_{U}$ increases for both mesh sizes. The NN errors are positioned below the $\left\lVert u-\pi_{h}(u)\right\rVert$ lines, demonstrating the improved NN accuracy without interpolations. Additionally, Fig. 7 validates the choice of $\gamma=10^{-2}$ for $k_{U}$ values ranging from 1 to 5, building on the previous subsections where $k_{U}\in\{2,3\}$.
@@
4.1.5. Discrete dual norms for the Poisson equation
In this experiment, we explore the impact of discrete dual norms in the training of unfitted FEINNs for the Poisson equation. We compare the performance of PINNs with both unfitted $\ell^{2}$- and $H^{-1}$-FEINNs. We use the same disk geometry as in Sect. 4.1.4 and set $k_{U}=3$. The background box $[0,1]^{2}$ is partitioned into $50\times 50$ quadrilaterals. To increase the difficulty of the problem, we use the following analytic solution, which features a sharp gradient around the centre of the disk:
@@ 4.1.5. Discrete dual norms for the Poisson equation
Since the solution is rougher than (11), the use of a discrete dual norm is instrumental in improving the convergence of the training process.
@@
As described in Sect. 2.4, we consider two linearised test FE spaces: the standard active space $V_{h}^{\rm{std}}$ consisting of all internal and cut elements, and the aggregated active space $V_{h}^{\rm{ag}}$ The discrete dual norm requires to compute the Gram matrix for the linearised test space with respect to the $H^{1}$-inner product. However, spectrally equivalent choices can also be considered. For example, in [Badia2024, Sect. 6.1.3], we demonstrated that a single cycle of GMG was an effective approximation of the inverse of the Gram matrix. Since small cut cells in $V_{h}^{\rm{std}}$ may cause the Gram matrix to become singular in the presence of rounding errors, additional ghost penalty terms (5) are required to ensure the Gram matrix is well-posed. Note that in the context of unfitted FEINNs, even though the trial space may be of higher order, $V_{h}^{\rm{std}}$ is always linear. Therefore, a first-order penalty term is sufficient to stabilise the Gram matrix, i.e., $k_{U}=1$ in (5). The implementation of unfitted FEINNs does not actually need to deal with high order derivatives in the case of CutFEM stabilisation, nor specialised discrete extensions required for high-order aggregated FE spaces [Badia2022c]. Following the recommendation in [Burman2015], we set $\gamma_{g}=0.1$ in this experiment. On the other hand, the $V_{h}^{\rm{ag}}$ space is designed to overcome stability issues by construction, thus no additional penalty terms are needed in the weak formulation.
@@
As shown earlier in Fig. 3, large Nitsche coefficients can slow down the training convergence, so we include both the suitable $\gamma=10^{-2}$ and a large $\gamma=10^{2}$ in the experiment to test how effectively $H^{-1}$-FEINNs handles unsuitable Nitsche coefficients. For clarity in the plots, in this experiment, we only consider the $H^{1}$ primal norm in the loss (9). However, other primal norms can also be considered. For example, in [Badia2025adaptive, Sect. 5.1.2], we showed that four different primal norms were effective for the Poisson equation.
@@
We plot the error histories for $\pi_{h}(u_{\mathcal{N}})$ and $u_{\mathcal{N}}$ for different combinations of test spaces for $\ell^{2}$-FEINNs and $H^{-1}$-FEINNs in Fig. 8. In the figure, we label $\ell^{2}$-FEINNs with the postfix “$V_{h}^{\rm{std}}$, $\ell^{2}$” and $H^{-1}$-FEINNs with the postfixes “$V_{h}^{\rm{std}}$, $H^{-1}$” or “$V_{h}^{\rm{ag}}$, $H^{-1}$”. We observe that for $\ell^{2}$-FEINNs, a large Nitsche coefficient also impedes the convergence of both $\pi_{h}(u_{\mathcal{N}})$ and $u_{\mathcal{N}}$ for this problem (with a solution featuring a sharp gradient around the disk centre). However, the use of dual discrete norms in $H^{-1}$-FEINNs overcomes this issue and accelerates the convergence significantly. For instance, with $\gamma=10^{2}$, the $L^{2}$ curve of $\pi_{h}(u_{\mathcal{N}})$ for $H^{-1}$-FEINNs drops to $10^{-5}$ at around 800 iterations, while $\ell^{2}$-FEINNs require more than 4,000 iterations to achieve the same error level. Besides, the use of a discrete dual norm in the loss function is also effective for the suitable Nitsche coefficient ($\gamma=10^{-2}$), as it does not not only accelerates the convergence of NNs but also enhances the NN solution accuracy. As shown in Fig. 8, the corresponding error curves decrease more rapidly than the $\ell^{2}$-FEINN ones. In addition, the NN error curves for $H^{-1}$-FEINNs in the second column of Fig. 8 drop below the $\ell^{2}$-FEINN ones shortly after training begins.
@@
It is clear from Fig. 8 that all the dual norm error curves for different $\gamma$ and test space combinations are very close. This implies that dual norm losses are robust to variations in $\gamma$ values and that the choice of test space has only a minor impact on their effectiveness. Since the Gram matrix requires neither additional stabilisation nor parameter tuning, we recommend using $V_{h}^{\rm{ag}}$ as the test space in practice. However, we reiterate that the unfitted FEINN framework only requires a first-order ghost penalty term to stabilise $V_{h}^{\rm{std}}$, which makes $V_{h}^{\rm{std}}$ also a practical choice.
In terms of computational cost, we note that the Gram matrix is symmetric positive definite and is built and factorised using sparse Cholesky factorisation only once before training. Since the problem is linear, this factorisation can be stored and reused throughout training. At each iteration, the inverse of the Gram matrix is applied to the residual vector by forward and backward substitutions using the sparse Cholesky factor, which makes the training computationally efficient.
[…]
In Fig. 9, we present NN error history versus training iterations and time in seconds. Let us first comment on the convergence curves of $\ell^{2}$-FEINNs and $H^{-1}$-FEINNs. The use of dual norms indeed improves computational efficiency. In the $L^{2}$ error plot on the left of Fig. 9, even with a suitable Nitsche coefficient ($\gamma=10^{-2}$), $\ell^{2}$-FEINNs require substantially more training time than $H^{-1}$-FEINNs to reach comparable accuracy. For instance, achieving the reference $\left\lVert u-\pi_{h}(u)\right\rVert$ error, as indicated by the red dashed lines, takes around 2,000 iterations and 3,000 seconds for $\ell^{2}$-FEINNs, compared to just 800 iterations and 1,000 seconds for $H^{-1}$-FEINNs. Another notable observation is that the overall computational cost of $\ell^{2}$- and $H^{-1}$-FEINNs is similar, with both methods finishing 3,000 iterations in around 4,000 seconds. This suggests that incorporating dual norms does not significantly increase the computational cost, while still accelerating convergence.
In the PINN experiments, we again tested different $\tau_{b}$ values and found that $\tau_{b}=10^{4}$ generally yielded the best results. Among the three methods, PINNs exhibit the slowest convergence in both time and iterations, requiring more than 2,700 iterations and 7,800 seconds to reach the reference $\left\lVert u-\pi_{h}(u)\right\rVert$ error; $H^{-1}$-FEINNs is noticeably the most robust method.
[…]